### v2

# “Medium” LMs of Code in the Era of LLMs: Lessons From StackOverflow

Abstract.
Large pre-trained neural language models have brought immense progress to both natural language processing and software engineering. Models in OpenAI’s GPT series now dwarf Google’s BERT and Meta’s RoBERTa, which previously set new benchmarks on a wide range of NLP applications. These models are trained on massive corpora of heterogeneous data from web crawls, which enables them to learn general language patterns and semantic relationships. However, the largest models are both expensive to train and deploy and are often closed-source, so we lack access to their data and design decisions. We argue that this trend towards large, general-purpose models should be complemented with single-purpose, more modestly sized pre-trained models. In this work, we take StackOverflow (SO) as a domain example in which large volumes of rich aligned code and text data is available. We adopt standard practices for pre-training large language models, including using a very large context size (2,048 tokens), batch size (0.5M tokens) and training set (27B tokens), coupled with a powerful toolkit (Megatron-LM), to train two models: SOBertBase, with 125 million parameters, and SOBertLarge with 762 million parameters, at a budget of just \$374 and \$1600 each. We compare the performance of our models with both the previous SOTA model trained on StackOverflow data exclusively (BERTOverflow, which relied on an approach tailored to natural language processing) as well general-purpose BERT models (BERTBase and BERTLarge) and OpenAI’s ChatGPT on four StackOverflow-specific downstream tasks - question quality prediction, closed question prediction, named entity recognition and obsoletion prediction (a new task we introduce). Not only do our models consistently outperform all baselines, the smaller model is often sufficient for strong results. Both models are released to the public [^1] https://huggingface.co/mmukh/SOBertLarge, [^2] https://huggingface.co/mmukh/SOBertBase. These results demonstrate that pre-training both extensively and properly on in-domain data can yield a powerful and affordable alternative to leveraging closed-source general-purpose models.
Large Language Models,Pre-trained Models, Discriminative models,Stack Overflow
booktitle: arxiv
1. Introduction
Neural Language Models (LMs) have brought about a significant transformation in the domains of Natural Language Processing (NLP) and Software Engineering (SE). In NLP, LMs are utilized for tasks including text classification (Sun et al., 2019), sentiment analysis (Song et al., 2020), and machine translation (Zhu et al., 2020). In SE, LMs treat code as a form of language and are used for tasks such as code completion (Liu et al., 2020), bug localization (Ciborowska and Damevski, 2022), and code summarization (Wang et al., 2020). Language models can be trained from scratch for a given task as long as they are provided with enough data, typically measured in millions of tokens. In recent years, however, it has been found more effective to pre-train models on large amounts of data to learn general patterns from text or code (Gururangan et al., 2020). Following this pre-training, smaller models such as BERT (Devlin et al., 2018) (with fewer than one billion parameters) are typically fine-tuned on downstream tasks while larger ones such as GPT-3 (Brown et al., 2020) (with many billions of parameters) can be prompted with a natural language question and some examples.
Figure 1. Model and dataset sizes of state-of-the-art LLMs.
Following the observation that language model performance scales predictably and steeply with compute budget (Kaplan et al., 2020), recent trends strongly favor larger models: models like GPT-3, with 175B parameters (Brown et al., 2020), and PALM (Chowdhery et al., 2022) with 540B parameters are trained with hundreds of billions of tokens using a fairly simple, text completion training signal. Commonly referred to as Large Language Models (LLMs) , these models gain remarkable capabilities on a host of language tasks when prompted with just a problem description and some examples. Figure 1 shows the relative scale in terms of both parameters and training data of various BERT models, including ours (in bold) compared to GPT-3 (note the log-scale on both axes). [^3] Where dataset size is only provided in tokens, we assume an average token size of 5 bytes. While GPT-3 is likely significantly smaller than state-of-the-art models in the GPT series, OpenAI has not released details on any more recent models. Yet the ease with which LLMs can be used for new tasks comes at a cost: LLMs are typically massive, requiring millions of dollars of compute to train (Hellendoorn and Sawant, 2021), making them nearly impossible to fine-tune on academic budgets. Most LLMs are also closed-source, so we have no access to their learned weights nor insight into their training data. Finally, as we will show in this work, their general-purpose training signal leaves them performing below-par on tasks where sufficient fine-tuning data is available.
The SE community has responded to the trend towards larger language models with code-specific LMs. For example, CodeBERT (Feng et al., 2020) (110M parameters) is based on BERT, a so-called encoder-only model trained to fill in masked-out tokens in an input. CodeT5 (Wang et al., 2021) (comes in 60M, 220M and 770 parameters) is an encoder-decoder model that encode inputs (mainly, entire functions) plus a task description and decodes (essentially, converts) these into outputs, such as labels, repaired functions. PolyCoder (Xu et al., 2022) (2.7B parameters), Codex(Chen et al., 2021) (12B parameters, not open-source) and CodeGen (Nijkamp et al., 2022) (16B parameters) are all decoder-only models trained to generate code in a left-to-right fashion. In this work, we add to this collection of models by training several BERT-style models on StackOverflow (SO). SO has the unique appeal of providing high-quality text (often, explanations) alongside code, making it a key resource for programmers. Generic (non-SO pre-trained) BERT models are frequently and successfully fine-tuned for downstream tasks involving SO data (Bleyl and Buxton, 2022; Biswas et al., 2020; Kou et al., 2023). Several papers even propose to pre-train such models on SO data, including BERTOverflow (Tabassum et al., 2020), seBERT (von der Mosel et al., 2021) and PTM4Tag (He et al., 2022), often finding that such models are even more powerful when fine-tuned for downstream tasks. However, these pre-training efforts tend to be based in NLP conventions, such as treating each sentence in a post as a separate input to the model. As such, they overlook key properties of StackOverflow posts and large language modeling best practices more generally, like the significant benefits of using very large inputs – BertOverflow limits inputs to 512 tokens, just like CodeT5; we use 2,048 in line with GPT-3 and corpus statistics shown in Figure 4. Unlike seBert (von der Mosel et al., 2021), we preserve the code within the text, as they contain information that is valuable to the overall comprehension of the content. Furthermore, we train models with both posts and their associated comments, which often add key information about the answer’s quality over time.
Figure 2. Framework outlining the key steps of our analysis
With this input representation, we train two BERT models on 27 billion tokens of SO data following best-practice guidelines for training large language models: SOBertBase, with 125M parameters, and SOBertLarge, with 762M parameters. While nontrivial in size, these models can be trained in the cloud at a budget of just \$374 and \$1600 respectively.
Figure 3. With this post (ID: 14569223) as an example, this image demonstrates the structure of a question and answer on SO, featuring a question title, question body, answer body, and comments. We also demonstrate which features we use for each of the following four downstream tasks.
We fine-tune both SOBert models as well as an array of baselines including two generic BERT models and BertOverflow, as well as ChatGPT queried via few-shot prompting, on four different tasks that measure StackOverflow question & answer understanding: question quality prediction (Annamoradnejad et al., 2022), closed question prediction(Kag, 2013), named entity recognition (Tabassum et al., 2020) and obsoletion prediction. The latter is a new task that we introduce focused on detecting whether answers contain outdated code, a common phenomenon on SO that can cause vulnerability and compatibility issues for those adopting the code without scrutiny (Fischer et al., 2017; Meng et al., 2018; Ragkhitwetsagul et al., 2019). We constructed this dataset by manually annotating about 1,000 SO answers. SOBert consistently, and often widely, outperforms all other compared models on all four tasks, with even the smaller model, SOBertBase frequently yielding state-of-the-art performance. These results highlight the significant promise of pre-training sub-billion parameter language models on in-domain data when such data is rich in detail and abundant. We release our models for public use.
In summary we make the following contributions:
  • We introduce two models, SOBertBase (125M parameters) and SOBertLarge (762M parameters), that were trained on StackOverflow answer and comment text using the Megatron Toolkit. The training process used multiple GPUs and cost a modest budget of \$374 and \$1600 for SOBertBase and SOBertLarge, respectively.
  • We extensively evaluate both SOBert models on four different tasks, which include question quality prediction, closed question prediction, named entity recognition, and obsoletion prediction.
  • Our evaluation involves comparing the performance of our models to two variants of BERT (BERTBase and BERTLarge), BERTOverflow, and ChatGPT (for three tasks). We demonstrate that SOBert outperforms all other models on all four downstream tasks, and sometimes even SOBertBase outperforms the other compared models in terms of performance.
2. Background
To provide context and prevent ambiguity for the features discussed later, we begin by defining some of the terms we will be using in this paper. We then delve into the background work that has been done in this field.
2.1. How StackOverflow Works
StackOverflow (SO) is a social Q&A platform. SO posts consist of several components, each of which serves a distinct purpose. The primary components of a post are the title, question body, answer body and comments. The title is the headline of the post, and it is meant to provide a brief summary of what the post is about. The question body is the main content of the post, where the user explains the problem they are facing or the question they want to ask. The answer body is where other users can provide their responses to the question asked in the post. Comments are additional remarks that users can add to a post. They can be used to ask for clarification, provide feedback, or suggest improvements to the post. Other components of a post include tags, which are keywords that describe the topic of the post, and votes, which allow users to express their opinion on the quality of the post. Figure 3 provides a visual representation of the various components that make up Stack Overflow (SO) posts. It namely illustrates the different elements, such as the title, question body, answer body and comments that users can interact with when creating or viewing a post on the platform. Users who ask questions are encouraged to both post text describing the problem and include other material such as their (minimal) incorrect code, screenshots, URLs etc. Questions should include between one and five ‘tags’, which identify features like the programming language, library used, or domain of the problem and serve to bring the question to the attention of experienced developers in the corresponding domains. Once a question is posted, developers may respond to with answers, one of which the question-asker may eventually mark as the accepted answer. It is not uncommon for multiple answers to be posted, and upvoted, for a single question; each usually differs in some aspect such as the library (version) used, assumptions about the underlying system & context, or degree of optimization. [^4] Indeed, popular questions frequently have at least one “meta-answer” benchmarking and comparing the solutions proposed in the other answers. Users have the ability to downvote answers or questions as well, to help reduce the visibility of poor quality content. It is not necessary for the accepted answer to have the most upvotes. Along with questions and answers, the website also features a commenting system that allows users to provide feedback or additional information on a particular question or answer. Comments on Stack Overflow can be used for a variety of purposes, such as clarifying a question, suggesting an alternative solution, or pointing out potential issues with an existing answer. They can also be used to ask for further information or to provide additional context that may be helpful in finding a solution to a particular problem.
2.2. Pre-trained Language Models
Language models generate text in a given context. Auto-regressive, or left-to-right, language models do so by repeatedly predicting the next token in an input conditioned on all previous tokens, while masked-language models predict a random subset (typically some 15%) of “masked” tokens in an entire input. Both can be pre-trained on very large datasets of text by converting each document into a series of training examples. Auto-regressive, or decoder-only, models tend to be useful for downstream tasks where new text is a desired output, as in ChatGPT, while masked-language model training assumes that the entire input document (minus zero or more masked-out tokens) is available while training which makes it easier to “embed” inputs in a single representation that may then be used for learning any given downstream task.
BERT (Bidirectional Encoder Representations from Transformers) is a type of pre-trained masked language model developed by Google introduced in (Devlin et al., 2018) that can be fine-tuned for various natural language processing tasks. It utilizes the Transformer architecture (Vaswani et al., 2017), which employs a self-attention mechanism to learn a contextual representation of text. It is trained using both the masked language modeling and next sentence prediction objectives on BooksCorpus (800M words) and Wikipedia (2,500M words). Of these two tasks, predicting masked words has proven particularly important as it teaches the model to represent each word in the context of its surrounding words, which helps it learn meaningful representations of the entire document as well. This makes BERT-like models versatile and effective in various downstream NLP tasks such as text classification (Sun et al., 2019; González-Carvajal and Garrido-Merchán, 2020) and named entity recognition (Labusch et al., 2019; Liu and Ritter, 2022).
Two well-trained BERT models were made available by Google. BERTBase-uncased (Devlin et al., 2018) is the smaller model, which has 12 layers of transformer blocks with an internal dimension of 768 for a total of 110 million parameters. The ”uncased” tag means that its input text has been converted to lowercase and all punctuation has been removed, which simplifies learning from a heterogeneous body of text. A second model, BERTLarge (Devlin et al., 2018), has 24 1,024-dimensional transformer layers for a total of 340 million parameters. We match our own Base model, SOBertBase, to the former, while adopting a model with a slightly larger hidden dimension (1,536) than BERTLarge but the same number of layers for SOBertLarge, yielding 762M parameters. We pre-train our models with the masked language modeling objective.
2.3. Modeling StackOverflow data
StackOverflow (SO) frequently releases ‘data dumps’ of publicly available content (questions, answers, comments, etc.). This, combined with its natural alignment of informative text and code, has made it a popular target for SE research. Numerous aspects of software engineering in have been studied on SO by researchers. For instance, one study scrutinized the interactions among developers on the platform (Wang et al., 2013), while another analyzed the themes and patterns in discussions on SO (Barua et al., 2014). Working with StackOverflow data is not without challenges. For one, the quality and structure of answers varies widely from post to post, with many answers being incomplete, containing errors, or lacking context. Another issue is the sheer volume of data on the platform, making it difficult to effectively navigate and identify high-quality task-relevant posts for specific study goals. To mitigate this, studies often focus on a narrow slice of the data, such as investigating the specific inquiries made by software engineers about Android testing on Stack Overflow (Villanes et al., 2017).
Post2Vec, as introduced in Xu et al.’s work (Xu et al., 2021), represents one of the pioneering methods for exploring the generation of distributed representations for Stack Overflow posts. Post2Vec achieved this through the utilization of Convolutional Neural Networks (CNNs) as the primary feature extraction mechanism. At the intersection of language models and SO data, there has been a lot of work in the space of fine-tuning BERT for various tasks such as emotion recognition (Bleyl and Buxton, 2022), sentiment analysis (Biswas et al., 2020), automated summarization (Kou et al., 2023) etc. that involve SO data. Several models have been pre-trained on SO data specifically. He et al.(He et al., 2022) build a tag recommendation framework PTM4TAG for Stack Overflow posts with different BERT-based models. The model seBERT (von der Mosel et al., 2021) is a 340 million parameter model trained on 62.8GB of textual data (they remove all code from the posts) from answer, posts and comments from SO along with more data from GitHub and Jira using the BERTLarge architecture. They use a cased-insensitive 30,522 token WordPiece vocabulary to tokenize their data and restrict inputs to a maximum length of 128 tokens – significantly shorter than most SO posts (Figure 4). Tabassum et al.(Tabassum et al., 2020) released a BERT model pre-trained on StackOverflow data. Their model, BERTOverflow, was pre-trained on 152 million sentences from SO using a case-sensitive 64,000 token WordPiece vocabulary. Training on individual sentences was a common practice in training BERT-like models, but recent work has found that language models perform substantially better when considering a document-level context (Kaplan et al., 2020), albeit at a substantial increase in compute cost. Our findings echo that observation. Tabassum et al. focus on the downstream task of Named Entity Recognition (NER, essentially tagging each word with its syntactic role), on which their model out-performs existing models such as BERTLarge as well as GloVe and ELMo models trained on the same SO data.
We selected BertOverflow as the main specialized baseline in our work. We did not include comparisons to the other models due to differences in model architecture and training data between these models and our proposed SOBert. Specifically, seBERT incorporates additional data sources beyond StackOverflow during pre-training, while PTM4Tag only models the Title, Description, and Code from StackOverflow posts. We omitted a comparison with Post2Vec because its performance has been substantially surpassed by Transformer-based approaches, e.g., as shown in PTM4Tag (He et al., 2022) on a tag prediction task. Instead, we include BERT-Base and BERT-Large as strong, general models, in line with our broader goal of demonstrating that domain-specific pre-trained models can outperform general models. And we include BertOverflow as a representative, capable domain-specific model to highlight how adopting LLM training practices allows for training significantly more powerful domain-specific models.
Compared to these efforts, we pre-train SOBert with 19 GB data presented as just 15 million samples where each sample contains an entire post and all its corresponding comments. We also include all code in each answer so that our model is bimodal in nature. Moreover, we use a SentencePiece tokenizer trained with Byte-Pair Encoding (Kudo and Richardson, 2018), which has the benefit over WordPiece of never labeling tokens as “unknown”. Additionally, SOBert is trained with a maximum sequence length of 2048 based on the empirical length distribution of StackOverflow posts (see Figure 4) and a relatively large batch size of 0.5M tokens (following Kaplan et al.(Kaplan et al., 2020)). Our model performs substantially better than both BertOverflow and BERT models trained on general corpora of text on all downstream tasks (see Section 5).
3. Approach
StackOverflow (SO) contains code in various languages with rich natural language information about the code. In this section we describe how we design SOBert, a BERT model trained on bimodal data from 15 million StackOverflow questions and answers.
3.1. Data and Preprocessing
We downloaded the SO data dump published on 7 March 2022 (SO, 2021) which includes posts from 2008-2022. The Stack Overflow data dump is a collection of structured data that includes all information publicly available on the website. The data is made available for download periodically and includes information such as user profiles, questions, answers, comments, tags, and votes. The data is provided in XML format and is compressed into a set of files that can be downloaded and processed using a variety of tools and programming languages. Each file corresponds to a specific type of data, and contains all the relevant information for that type. We loaded these files into SQL database tables and specifically worked with the ‘Posts’ and ‘Comments’ tables. Using the entire corpus of Stack Overflow posts would have introduced a significant amount of low quality, unengaged content that may not be representative of posts software engineers typically rely on. We thus make a design choice to prioritize quality content that demonstrated some level of community engagement by filtering the answer posts to extract only those that meet the criteria of having a minimum of one up-vote and at least one accompanying comment. We then use this filter and extract both answer posts along with all their associated comments. This yielded 15 million answers and 39.5 million comments (median 2, mean 2.68 comments per post).
3.2. Data decisions
Devlin et al. (Devlin et al., 2018) and Tabassum et al. (Tabassum et al., 2020) tokenize the datasets for BERT and BERTOverflow respectively using a WordPiece vocabulary with 30,522 tokens. We preprocess the data for SOBert using a 50,000 token Byte-Pair Encoded (BPE) vocabulary, trained using SentencePiece (Kudo and Richardson, 2018). BPE sub-tokenizes words by greedily finding the largest tokens in its vocabulary that collectively reconstruct the word. While common words such as “the” tend to appear as a single token in the vocabulary, rare words may be broken up into smaller pieces. Since BPE uses a large vocabulary of single characters as its base vocabulary, it avoids treating almost any token as “unknown” or out-of-vocabulary, in contrast to WordPiece, which tends to improve results on downstream tasks (Karampatsis et al., 2020). We trained our tokenizer using a randomly selected subset of 10% of the entire dataset (ca. 2GB) due to memory limitations. The input text is first cleaned following standard data cleaning procedures such as first replacing URLs and email addresses with generic [URL] and [EMAIL] tokens to abstract away identifiable information.
We make sure to keep blocks of code, demarcated with <code> </code> tags, in the input data to preserve the complete structure of each post. This ensures that any code snippets or programming language syntax are not altered or misrepresented during processing or analysis. The preprocessed text is then input into the SentencePiece tokenizer for subword tokenization. A reserved token, <RS>, is placed between the answer and each of the comments to signal to the model where each comment begins. This tokenized representation is then passed through a stack of Transformer layers, described next.
3.3. Model Design
Figure 4. Histogram showing length buckets of post+comments samples
For the BERT base model the maximum length of the input sequence for the model is 512 tokens and each token is represented with a 768 dimensional vector. The choice of 2048 as the maximum length of the input sequence for SOBert was informed by an analysis of the text lengths of combined posts + comments that we use as inputs to our approach, shown in Figure 4. This shows that rather few posts are no more than 512 tokens long, which has conventionally been used as the maximum input length (Feng et al., 2020; Wang et al., 2021). A slight majority of the posts contain up to 1024 tokens but many more posts span between 1024-2048 tokens. As the cost of encoding long inputs grows quadratically in attention-based models, we choose to set the maximum length of inputs in our study to 2048 to balance training cost and fidelity to the task domain.
Batch size refers to the number of training examples that are processed together in a single iteration during training. Larger batch sizes tend to lead to better language model performance (Kaplan et al., 2020) by providing more stable gradients for each model update, but they also require more memory and computing resources. We use gradient accumulation to simulate large batch sizes on our limited hardware, allowing us to process the training data in batches of 0.5 million tokens for both models. DeepMind proposed a series of “scaling laws” that can be used to determine how to best balance a given compute budget between training larger models and providing models more data (Hoffmann et al., 2022). Their conclusion was that models should be trained with ca. 20 times more tokens than they have parameters. While models trained this way won’t have converged (training longer would continue to yield better performance), the empirical observation is that this minimizes loss for a given compute budget. For instance, if one had the budget to train a 1 billion parameter model with 50B tokens, it would instead be better to train a significantly larger (e.g., 1.5B parameter) model with 30B tokens. The latter would achieve a lower loss in the same training time. This “rule” thus provides a useful lower bound for the training budget that should be allocated to a given model. We determined that, with our compute budget (see Section 3.4), the largest model that we would be able to train to this “optimum” within ca. two weeks time contained ca. 760M parameters. We therefore aimed to train this model with at least 15B tokens. Our actual training resources enabled us to provide nearly twice as many tokens as this minimum, which naturally improved its performance (Figure 5). [^5] We follow the conventional GPT-2 architecture sizes here. The typical “next-size” model would contain 1.4B parameters, which we estimate we could only train with 10 - 15B tokens, falling well short of the compute-optimum.
We train our models using the Megatron-LM toolkit (Shoeybi et al., 2019). This toolkit was introduced for efficient parallel model training of LLMs with up to 8.3B parameters. It provides powerful support for multi-GPU training with optimized kernels for many components. It also incorporates many architectural best-practices. Shoeybi et al.(Shoeybi et al., 2019) showed that rearranging the order of the layer normalization and the residual connections is critical to enabling the scaling of the BERT-style models beyond 336M parameters. We use the same architecture configuration for SOBert.
Figure 5. Training and Validation Loss comparison of SOBBase and SOBLarge.
3.4. Training details
The training was performed on four NVIDIA Quadro RTX 8000 GPUs, each with 48GB of memory, hosted in an on-premise server. To calculate the cost of training our models for those without access to this hardware, we base our estimates on quotes for NVIDIA’s A6000 GPU series, as these are currently widely available via cloud computing environments and have the same memory capacity as our GPUs. Nevertheless, A6000 GPUs outperform the Quadro RTX 8000 by ca. 80% (gpu, 2023b). We thus take the GPU hours required to train each of our models and multiply them by the standard rate of an A6000 in the cloud, which is approximately \$1 per hour (gpu, 2023a), and by the ratio 1/1.8 to compensate for the performance disparity. Training the SOBertLarge model required ca. 30 days on four GPUs for a total of ca. 2,880 GPU hours, which is equivalent to \$1600. The SOBertBase model required just 7 days ( ca. 672 GPU hours) to train, which amounts to \$374.
Figure 5 shows the comparison between the training and validation loss curves of SOBertBase and SOBertLarge. Both SOBertBase and SOBertLarge were trained for 100,000 steps at which point they had seen around 54 billion tokens.
4. Evaluation
In this section we define the metrics we use to evaluate and compare the models and describe each of the four StackOverflow (SO) related downstream tasks, most of which come from prior research or Kaggle competitions. The four tasks we consider are question quality prediction (Annamoradnejad et al., 2022), closed question prediction(Kag, 2013), named entity recognition (NER) (Tabassum et al., 2020) and a novel task we formulate called obsoletion detection. We try to cover different types of downstream tasks with varying amounts of data available for fine-tuning. The above tasks, described in detail below, include a multi-class classification problem with about 60k samples, a multi-class classification problem with 140k samples, a token-level tagging task with 15k samples and a binary class classification with just 942 samples.
Relevance of tasks to Software Engineering (SE) community: We have selected these four tasks based on their significance within the software engineering (SE) community and the availability of data for effective model performance comparisons. The first task involves predicting the quality of questions on platforms like StackOverflow. By identifying low-quality questions, this task contributes to a more efficient and informative knowledge-sharing environment. The second task predicts question closures, assisting moderators in managing platform content and ensuring its quality. Named Entity Recognition (NER), the third task, extracts crucial technical entities from text, benefiting various SE tasks such as code search by linking APIs and libraries to documentation. Finally, obsoletion detection, our fourth task, plays a vital role in highlighting outdated code snippets, promoting code health and security by preventing the use of vulnerable, deprecated code. These tasks collectively empower the development of improved tools for software developers, ultimately contributing to a more efficient and secure SE environment.
Our evaluation of the models involves a comparative analysis with two general-purpose BERT models, namely BERTBase (Devlin et al., 2018) and BERTLarge (Devlin et al., 2018). In addition, we also assess the performance of the models against a domain-specific model called BERTOverflow (Tabassum et al., 2020) as well as with ChatGPT (cha, 2023), although we do not fine-tune the latter. In all cases, each model is trained with the same hyper-parameters, typically involving a batch size of 32 samples and a learning rate of 1e-5. Small variations from task to task, mainly in the input length, are noted below. We note that, although our model is capable of handling longer sequences, we set the maximum sequence length to be the same as that of the baseline models for all our experiments.
SOBert was pre-trained using the Masked Language Modeling (MLM) objective. This means that during its pre-training phase, it was exposed to a wide range of text from Stack Overflow, which may or may not include some answers that appear in the downstream tasks. However, it did not have access to the labels for any of the downstream tasks. This reduces the risk of confounding effects related to data leakage in downstream tasks.
4.1. Metrics
Two of our tasks (NER and Obsoletion) have imbalanced datasets. An imbalanced dataset is one in which the distribution of classes is not uniform, meaning that one class may be over-represented or under-represented in comparison to the other classes. If the model always predicts “not-obsolete” or “O” the recall will be high, while a model that nearly never predicts “obsolete” or any of the other NER tags except when highly certain will achieve high precision. Metrics such as accuracy also tend to overvalue the performance on the majority class in imbalanced datasets. Class-weighted accuracy adjusts the contribution of each class to the overall accuracy score based on the inverse of the class frequency, giving more weight to the underrepresented classes. This offers a better indicator of a model’s true performance. The formula for weighted accuracy is:
$Acc_{W}=\sum_{i=1}^{N}\frac{n}{n_{i}}*Acc_{i}$
Where $n_{i}$ is the number of instances belonging to class $i$, $n$ is the total number of instances in the dataset, $N$ is the total number of classes, and $Acc_{i}$ is the accuracy of the model on class $i$.
Similarly, weighted recall is a performance metric used in classification tasks to measure the ability of a model to correctly identify all positive samples, taking into account the class imbalance. Recall is a measure of how well the model can identify positive samples from the total number of positive samples in the dataset. The formula for weighted recall incorporates class-weights as follows:
$Recall_{W}=\frac{\sum_{i=1}^{N}w_{i}*TruePositives_{i}}{\sum_{i=1}^{N}w_{i}*ActualPositives_{i}}$
Where $n$ is the number of classes, $w_{i}$ is the weight assigned to each class $i$, $\text{TruePositives}_{i}$ is the number of correctly predicted positive samples for class $i$, and $\text{ActualPositives}_{i}$ is the total number of actual positive samples for class $i$. The class weights are typically selected according to their inverse frequency, as in the weighted accuracy formula.
Neither $Precision$ nor $Recall$ fully captures a model’s performance. $F1-score$ takes both $Precision$ and $Recall$ into account and is their harmonic mean. Its class-weighted formula is:
$F1_{W}=\frac{\sum_{i=1}^{N}w_{i}*F1_{i}*ActualPositives_{i}}{\sum_{i=1}^{N}w_{i}*ActualPositives_{i}}$
In which $F1_{i}$ is the F1 score for class $i$.
4.2. Task 1: Named entity recognition (NER)
Tabassum et al.(Tabassum et al., 2020) release a dataset for the extraction of software-related named entities from StackOverflow. It spans roughly 10 years (from September 2008 to March 2018) of question-answer threads in which each token in each sentence is manually annotated with close to 50 types of entities. The StackOverflow NER corpus has 9,352 train, 2,942 development and 3,115 test sentences. We fine-tuned all the models with a batch size of 16, maximum sequence length of 128 (based on the maximum length of the samples in the dataset) and learning rate of 1e-4.666Hyper-parameters are particularly different here because this task involves one label per token, so a batch of 16 samples has far more labels than in the other, sequence-level prediction tasks. We use a weighted cross entropy loss function as the dataset is imbalanced with many more ‘O’ (no label) tags than others. We were able to replicate the results of (Tabassum et al., 2020) within a 5%-10% margin for all reported metrics with this configuration on both their dev and test sets. The results of this task are discussed in the next section.
4.3. Task 2: Quality Tag prediction
We use another Kaggle competition dataset for predicting question quality on SO (Annamoradnejad et al., 2022) for this downstream task. This dataset features 60,000 SO questions from 2016-2020 classified into three categories based on their quality:
  • HQ: High-quality posts without a single edit.
  • LQ_EDIT: Low-quality posts with a negative score, and multiple community edits. However, they still remain open after those changes.
  • LQ_CLOSE: Low-quality posts that were closed by the community without a single edit.
We use a batch size of 32, learning rate of 1e-5 and set the maximum sequence length as 512 for this task. The results of this task are discussed in the next section.
4.4. Task 3: Closed question prediction
We choose a Kaggle competition dataset for predicting closed questions on SO (Kag, 2013) for this downstream task. The provided dataset contains 140,272 samples. The competition motivates this task by saying that SO is a widely-used platform that provides programmers with high-quality answers to their programming questions on a daily basis. At the time this competition was published, SO received over 6,000 questions every weekday and an automatic content moderation would help maintain post quality. One aspect of moderation involves closing questions that are deemed inappropriate for the platform. At the time this competition was published, around 6% of all new questions ended up being closed for various reasons. The purpose of the competition is to analyze said quality and predict when questions should be closed for effective moderation. They excluded the ‘exact duplicate’ closed reason from the dataset as it depends on previous questions. The competition requires the development of a classifier that can predict whether a question will be closed or not, based on the submitted question and the reason for closure. The categories of classification are ‘open’, ‘not a real question’, ‘off topic’, ‘not constructive’ and ‘too localized’.
The competition provides a plethora of features from user data, however for simplicity we use the question title and question body for the classification task at hand. We use a batch size of 32, learning rate of 1e-5 and set the maximum sequence length as 512 for this task. The results of this task are discussed in the next section.
4.5. Task 4: Obsolete post detection (New)
Figure 6. Answer 4601538 marked as obsolete on StackOverflow
Given the speed at which technology advances, answers on StackOverflow often become outdated. Outdated answers are any of those that propose solutions that are no longer state-of-the-art. They may offer solutions that are less performant or precise than what is now available, use now-deprecated APIs and libraries (be obsolete), and in some cases propose solutions that create security concerns. For instance, there have been more than 25 versions of TensorFlow (tf, 2021) since its release in 2015, including a major change to its entire compute structure (version 2). This library is also a highly popular subject of questions. Many of those are now years old and involve classes, methods, and parameters that have since been renamed. More generally, over 24 million answers contain code snippets, which software engineers often hope to reuse. When such snippets are obsolete, that reuse may severely reduce the quality of their code.
One might hope that such “outdatedness” is easily discovered by e.g. looking at the date posted, or comments pointing out such issues. However, date is not always a good indicator of obsolence (one type of outdatedness); Zhang et al. (Zhang et al., 2019) found that more than half of obsolete answers were already obsolete at the time of posting. Indeed, ‘year of posting’ only poorly explained outdatedness in our analysis as well. Comments are a useful resource: many posts have comments to indicate obsoletion. However, commenting or editing answers is time-consuming, so many posts with obsolete content go unmarked. Downvoting obsolete answers is ineffective, as posts that were once correct and popular require many downvotes to yield a negative score. SO’s maintainers are aware of this problem and have recently proposed explicitly marking outdated answers as such as shown in Figure 6. However, there are only 9 such answers marked as obsolete by SO at the time this paper was written. We thus propose to fine-tune SOBert to be able to predict these obsolete answers based on a manually curated dataset, which we describe next.
4.5.1. Raw Data Collection
We curate an initial dataset of obsolete answer candidates based on three heuristics described below, which are selected to provide the model with a wide range of patterns to learn from. In all cases, we focused only on answers with at least some community support, filtering out answers which had less than 1 upvote to reduce noise from potentially bad solutions.
Comments With Keywords
Users often leave comments on a post or answer to indicate obsoletion. We selected answer threads on the basis of the selection criteria from (Zhang et al., 2019). Comments were chosen which had one of the four keywords: deprecated, outdated, obsolete or out of date, while the corresponding question posts did not have these keywords in them. We found 85,586 such samples of comments.
Answers Edited After Comment
Often, though far from universally, when comments point out issues with an existing answer, that answer is edited to resolve that concern. Yet it is also not uncommon for answers to be edited even without such prompting, sometimes for cosmetic reasons, often soon after posting. Other edits are more substantial and often dramatically revise and extend an answer. To prioritize discovering the latter kind, we chose answers that were updated after a comment was posted and the Levenshtein distance between the code before and after edits was at least 100, indicating a significant change to the answer. We also filter out answers which have been included while considering the comments with keywords to prevent these answers from being repeated in our dataset. We find 388,809 samples of answers edited substantially following a comment.
Answers Added Late
Often, questions continue to draw attention for years after posting from developers facing similar issues, but the originally accepted answer ceases to be relevant. When this happens, the simple solution would be to simply edit the existing answer, but for various reasons [^7] Some related to SO’s “reputation” system, which prevents all-but highly experienced users from editing others’ answers. this is not always the path taken. Rather, we found many cases where a new answer was posted long after the original/accepted one, which usually references the previous answers in some way. These may serve various purposes, such as showing how a new feature can be used for a better solution than what was previously posted. We included such answers that were posted 1.5 years or more after the initial question was posted, where the answer had more than one upvote and referred to another answer on the same question thread. The check for reference to another answer is heuristic and keyword based, where we look for the keywords ’s answers, answer by, accepted answer and other answer. We found 19,371 such samples.
4.5.2. Data Annotation
After applying our heuristic selection criteria described above we obtain some 493K samples from the total database. We manually annotate 1000 random samples out of these, sampling equally from the three categories. The goal of the annotation process is to produce samples consisting solely of the text in an answer or comment on the answer along with a classification of obsolete or otherwise. Answers or comments that either include or refer to obsolete code in any way (e.g., comments pointing out obsolete code in the answer or answers edited to point out they contain outdated solutions) are classified as obsolete, regardless of whether newer techniques or alternative solutions are also provided in the same. This way, models trained on this dataset can be used to highlight posts that should be marked as obsolete using SO’s new system based on existing community input.
The annotation protocol was developed by having two professional programmers annotate a pilot set of answers independently. After completion of the pilot annotation, the differences and difficult-to-annotate cases were discussed and details of such cases were documented. This protocol was improved iteratively by both professionals on other pilot sample sets until consensus was reached about the protocol. The data annotation process relied on the protocol and a set of 10 annotated samples that served as guidelines.
The pilot annotation sets were then discarded and the same samples were re-annotated using the finalized protocol. The inter-annotator agreement of the annotations using the finalized protocol has a Cohen’s kappa of 0.63, which indicates that the agreement level is substantial (Landis and Koch, 1977).
This yielded many non-obsolete examples – while many comments and new answers contain the keywords we relied on, and many answers are heavily edited over time, these often signified spurious concerns, so the task for the models is nontrivial. We fine-tuned all the models with a batch size of 32, maximum sequence length of 512 (this being the largest sequence length supported by the baseline models) using a learning rate of 1e-4. We use a weighted cross entropy loss function as the dataset is imbalanced with many more non-obsolete samples than obsolete ones.
Task Model Weighted Accuracy Weighted Precision Recall F1-Score
[Uncaptioned image] NER
BertOverflow
0.69
0.67
0.68
BBUBase
0.80
0.46
0.55
BBULarge
0.56
0.83
0.56
0.63
SOBertBase
0.88 0.93
0.88
0.90
SOBertLarge
0.83
0.86
0.93
0.83
SoftNER
0.78
0.80
0.79
[Uncaptioned image] Quality
BertOverflow
0.72
0.72
0.72
0.72
BBUBase
0.82
0.83
0.82
0.82
BBULarge
0.71
0.72
0.71
0.71
SOBertBase
0.86
0.86
0.86
0.86
SOBertLarge
0.92 0.92 0.92 0.92
[Uncaptioned image] Closed Question
BertOverflow
0.53
0.49
0.53
0.41
BBUBase
0.60
0.56
0.60
0.56
BBULarge
0.53
0.48
0.53
0.43
SOBertBase
0.63 0.62 0.63 0.61
SOBertLarge
0.63
0.61
0.63 0.61
[Uncaptioned image] Obsoletion
BertOverflow
0.17
0.98
0.17
0.26
BBUBase
0.72
0.97
0.72
0.82
BBULarge
0.59
0.97
0.59
0.72
SOBertBase
0.87
0.97
0.87
0.91
SOBertLarge
0.90 0.97 0.90 0.93
Table 1. Performance Metrics for the Models Across the 4 Tasks - Named Entity Recognition (NER), Quality Prediction, Closed Question Prediction, and Obsoletion Detection
5. Results and Discussion
Figure 7. Comparison of F1-score performance of various models on 4 down-stream tasks: (1) NER, (2) quality tag prediction, (3) closed question prediction, and (4) obsolete prediction. The models evaluated include BERT (base and large), SOBert (base and large), BERTOverflow. We include SoftNER for the NER task. We also include ChatGPT@0 and ChatGPT@3, which consist of zero-shot and 3-shot prompting ChatGPT respectively, for closed question prediction, quality tag prediction and obsoletion detection on 250 random samples from the test set. SOBert outperforms the other models on all four tasks.
We fine-tune the SOBert models and the three baseline BERT models on the four downstream tasks described in 4. Figure 7 shows the F1-scores attained by the different models juxtaposed with their respective sizes (note the log-scaling and broken x-axis).
In the NER task, we utilized the test set provided by the authors and observed a significant improvement of +22 points in the F1-score when comparing SOBertBase to BERTOverflow as shown in Table 1. For this task, we did not compute the weighted accuracy of models other than the SOBert models, since the original paper (Tabassum et al., 2020) did not incorporate it in their evaluation metrics. We found that SOBertLarge had a greater improvement of +19 in F1-score over BERTLarge, with SOBertBase coming in close behind with an increase of +18 in F1-score over BERTLarge. Both SOBert models had an increase of about +28 over BERTBase. SOBertLarge also had higher accuracy, precision, and recall values of 0.83, 0.93, and 0.83, respectively, compared to BERTLarge’s scores of 0.56, 0.83, and 0.56. Additionally, when compared to the values reported in the paper for SoftNER, which incorporates a context-independent code token classifier with corpus-level features on top of BERTOverflow to further increase performance, we found that both SOBert models had an improvement of ca. +11 in F1-score over SoftNER.
In the second task quality tag prediction (results in Table 1), we found the SOBertLarge model to be particularly powerful, yielding a +20 point F1-score improvement compared to BERTOverflow. Our smaller, SOBertBase also improved over BERTOverflow’s F1 score by +14 points and by +15 points over BERTLarge.
In the third task (results in Table 1), which involves closed question prediction, we observed an increase of +20 in F1-score between both SOBert models and BERTOverflow. This marks the second task, after NER where our Large model is slightly outperformed by the smaller one – we discuss this phenomenon below. The F1-score for closed question prediction was relatively low for all the models in this task, with the highest F1-score achieved by SOBertBase and SOBertLarge at 0.61. While low, these scores are similar to the results reported in recent studies on the same task (Roy and Singh, 2020). The dataset used in this task contains several features, such as CreationDate, Tags, Reputation, Title, BodyMarkdown, etc. Roy and Singh(Roy and Singh, 2020) have explored the use of feature selection techniques and chose to use BodyMarkdown. We used both the Title and BodyMarkdown.
In the fourth task (results in Table 1) which involves obsoletion detection, we see large gains on the F1-score between SOBert and all the other baselines models. The F1 score of BERTOverflow is significantly lower than that of all other baselines. In 5.2 we delve into potential causes for this behavior and explore possible explanations.
5.1. Relation to BERT
In several experiments, the generic BERT models outperform BERT-Overflow, despite the latter using a very similar architecture and training on StackOverflow data only. In two cases, quality tag and closed question prediction, they even came close to rivaling one of our SOBert models. Their strong performance can likely be attributed by a combination of two causes. Compared to BertOverflow, these BERT models trained on significantly more data (16GB vs. 11GB), which would have helped them learn the patterns in language more effectively. Secondly, the two downstream datasets where they perform best are relatively large, containing 60K and 140K samples respectively. Fine-tuning on a larger dataset allows these models to better overcome the distribution-shift compared to their pre-training setup.
5.2. Relation to BERTOverflow
For the final task of obsoletion detection, we find the F1-score increased by a large margin between both SOBert models and BERTOverflow. Although BERTOverflow converges the fastest during fine-tuning between all 5 models compared for this task, we see that it struggles with this classification task. There could be various factors contributing to unsatisfactory performance, which may include some design decisions taken during pre-training the model such as a small context size and the structure of training data. The SO data BERTOverflow was pre-trained on processed individual sentences in isolation which has lead to a loss of contextual information from the original SO corpus. It is also possible that BERTOverflow was inadequately pre-trained, thus rendering it unable to generalize well to newer data and tasks.
5.3. Relation to model size
In this study, we noticed that both SOBert sizes perform better than both BERT sizes in all of the four tasks we examined. However, we did not find a consistent trend of the larger models (of either kind) outperforming their smaller counterparts. For example, in closed question prediction and obsoletion detection, SOBertBase outperformed SOBertLarge in terms of F1-score. Similarly, BERTLarge did not perform as well as BERTBase in quality tag prediction. Out of the four tasks, the two large models both performed better in exactly half. While this runs counter to the wisdom that larger language models are better (Kaplan et al., 2020), we note that this notion typically applies to models that are not fine-tuned but rather prompted (queried). Fine-tuning adds the noteworthy constraint that larger models represent their inputs in more hidden dimensions, which means that the task-specific prediction layer(s) also require more parameters. For instance, predicting a 1-dimensional label from a 768-dimensional hidden state (for the Base models) requires training half as many new parameters as transforming the 1,536-dimensional states from the Large models. Tuning more parameters also increase the threat of overfitting as models quickly learn to memorize their training inputs on small datasets.
A noteworthy observation from the training process was that, while loss continued to decline (5), downstream task performance saturated around halfway through training, at approximately 50,000 steps, for both SOBertBase and SOBertLarge. While language models generally continue to improve on end tasks as the training loss decreases, this trend does not necessarily hold for models that involve a separate task-specific fine-tuning phase. The additional fine-tuning step appears to enable BERT architectures to reach peak downstream performance well before full convergence of the base pre-trained model.
This highlights that in-domain BERT models need not be exhaustively pre-trained in order to realize most of the gains from domain-specific optimization. Had we halted training at 50,000 steps, we would have incurred only about half of the total compute cost, while likely retaining the bulk of the performance gains over general-domain BERTs demonstrated in this work. This insight could impart significant cost savings when pre-training models for new domains.
5.4. Relation to ChatGPT
Finally, we compared the results from SOBertBase and SOBertLarge with ChatGPT on 3 downstream tasks – closed question prediction, question quality prediction and obsoletion. While it would be great to be able to test on more samples, due to limits of testing ChatGPT, we mainly evaluate ChatGPT on a small part of the testing set for each of these tasks. Specifically, we chose 250 random samples from the test data (using test+validation for obsoletion detection, the smallest dataset). We use the GPT3.5 API to get predictions from ChatGPT. We get predictions from both SOBertBase and SOBertLarge on the same 250 samples. We compare this with responses from ChatGPT in 2 settings - with only an instruction and no examples (0-shot prompting) and with three examples drawn from the training data (3-shot prompting). The F1-score results for each of the tasks are shown in Table 2. These results indicate that although ChatGPT can solve many NLP problems quite well, it doesn’t fare too well on tasks in this domain.
A point to consider is that our downstream tasks are essentially classification tasks, and ChatGPT is a decoder-only pre-trained language model as opposed to SOBert which is an encoder-only model. BERT-like models are built upon a bidirectional Masked Language Modeling (MLM) objective that emphasizes the encoding of contextual information. Decoder-only models instead learn to derive representations that are primarily useful for text generation, which prioritizes knowledge from recent tokens. This, combined with its lack of domain-specific fine-tuning (in exchange for much more general capabilities) could explain the lower performance for ChatGPT on these tasks.
Task ChatGPT@0 ChatGPT@3 SOBertBase SOBertLarge
[Uncaptioned image]   Quality
0.39
0.44
0.9
0.89
[Uncaptioned image]   Closed Question
0.16
0.17
0.65
0.64
[Uncaptioned image] Obsoletion
0.56
0.52
0.90
0.88
Table 2. Comparison of task accuracy for SOBertBase, SOBertLarge, ChatGPT@0 and ChatGPT@3.
We started this work by arguing that, while many tasks generally benefit from general purpose LLMs, they do not necessarily guarantee cutting-edge performance for tasks where rich domain-specific data is available. The superior results of even our smallest model, SOBertBase, over both ChatGPT and several general-purpose BERT models on our downstream tasks shows that model size is not the only important factor for performance. Our substantial improvement over BertOverflow further highlights that it is crucial to carefully evaluate the requirements, constraints and data available for a given problem and use an appropriate training design to address the specific challenges at hand.
5.5. Limitations
While our assessment of SOBert includes a comprehensive analysis of four downstream tasks from well-established sources, it is important to recognize that the applicability of our findings to different tasks may differ. We also did not include comparisons with GPT-4 in our study due to its unavailability during our experiments. Additionally, the nature of API access from OpenAI implies that the performance of the language model can fluctuate over time due to model updates without prior notification, potentially affecting its performance. We were also restricted by API limitations for the number of samples we could prompt for and the number of examples provided in the prompt. However, the fact that the LLM is much less accurate in each task implies that it is unlikely to be able to bridge the absence of task specific training with examples alone.
6. Conclusion
We present SOBertBase and SOBertLarge, 125M and 762M parameter models respectively trained exclusively on StackOverflow data. We discuss a number of design decisions inspired by both the training of large language models and the nature of StackOverflow posts that enable us to train our models to a high quality. Among others, we create large samples consisting of entire StackOverflow posts and all their associated comments; we train with large batch sizes using a toolkit that enables efficient parallel training, and train our models extensively. With these configurations, the models we trained could be reproduced with cloud computing resources at a budget of just \$374 and \$1600. We then fine-tuned both our models and an array of BERT baselines on four different downstream tasks: question quality prediction, closed question prediction, named entity recognition and a novel task, obsoletion detection. We compare all our models, as well as ChatGPT (on all but one task) and find that our models consistently achieve top performance in all downstream tasks, with the smaller model regularly performing best. Overall, we underscore the importance of considering the specific needs of a task and the available data when choosing an appropriate model training approach, and provide our guidelines for doing so effectively. Our results highlight that there is ample room for smaller, domain-specific models to outperform general-purpose LLMs at a low budget.
7. Data Availability
Our supplemental materials are available on Figshare (fig, 2023). That includes the newly introduced obsolete answers dataset, our annotation protocol, and both SOBertBase and SOBertLarge models.